assume that a milliliter of water contains 20 drops. How long, in hours, will it take you to count the number of drops of water in 1.0gal of water at a counting rate of 10 drops/s?
1
Expert's answer
2019-08-26T05:46:00-0400

Let's take that 1.0 gal = 3.785 l

Then it contains 3.785 × 20 = 75700 drops

With a counting rate 10 drops/s it would take 75700 / 10 = 7570 s or 7570 / 3600 = 2.1 hours to count the number of drops.